OpenStudy (anonymous):

\[(1+xy)y dx + x(1-xy)dy = 0\] Get the general solution.

OpenStudy (anonymous):

Hello Guest, I hope you won't leave me (:forever alone:)... and then finally maybe somehow post the solution.

OpenStudy (anonymous):

the IF is 1/(2x^2y^2) after multiplying with IF one get ((1/2x^2y) +1/(2x))dx +((1/2xy^2) -1/(2y))dy =0 thus the equation becomes exact now after integ we have -1/(2xy) +1/2 lnx -1/2 lny = c (u can adjust c as per req) hope this solves ur problem suggest if i am not correct

OpenStudy (anonymous):

You are right, thanks! :-)

OpenStudy (anonymous):

may i know which book ur following for ur grads